Six coins are tossed 6,400 times. Using Poisson distribution, find the
approximate probability of getting six heads two times.

Grade:12

1 Answers

Nishant Vora IIT Patna
7 years ago
Let the coin be unbaised coin

Therefore the probability of getting 6 heads with six coins = (1/2)^6 = 1/64 = p (say)

Then np = 6400 * 1/64 = 100=m (say)

Then according to poissons law

P(X=x) =\frac{e^{-m}m^{x}}{x!}

=e^{-100}\frac{100^2}{2!}

=5000e^{-100}